Given an arrival process with λ= 5.0, what is the probability that an arrival occurs after t-2 tim e units? The probability is □ Round to four decimal placesas needed.)
Answer #1

(given rate) P(X < x) = 1-e(-5x) for x > 0 → P(X > x) = e-k -Ax P(X > 2)-exp(-12) e-10 0.00004540 0.000045 P(X > 2) 0.0000